OpenStudy (anonymous):

A 3.0kg block of wood is pushed along the horizontal flat surface with a force of 50.0 N. If the friction force is 10N, what is the acceleration of the box?

OpenStudy (anonymous):

So what does the free body diagram for this look like? ^_^

OpenStudy (anonymous):

 \[ \sum F_y = N-mg=0 \\ \hspace{100px} N=mg\] \[ \sum F_x = F_A - F_f = ma \\ \ \\ \quad \quad F_A = 50N \\ \quad \quad F_f = 10N \\ \quad \quad m=3 \ kg\] \[a= \frac{F_A - F_f}{m}\]